Find the inverse of f(x)=x^3 + 2
Answer by RAY100(1637)

f(x) = x^3 +2
.
y= x^3 +2
.
for inverse let y=x,,,and x=y
.
x=y^3 +2
.
this can be rearranged as required,perhaps,,,y = (x-2)^1/3